• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name mercado

    Log entry time 10:56:19 on August28,2009

    Entry number 286551

    keyword=Calibrating BCMs for 100uA

    In preparation for 100uA beam, I moved around some of the BCMs in order to get 
    healthy signals from our 16bit ADCs in the Counting House.
    BCMs 2 and 5 were moved:
    BCM2 was in 22_1. Moved to 15_0.
    BCM5 was in  2_1. Moved to 22_1.
    All boards should be run at LO gain.
    Board 22 MUST be on HI gain (~40000 signal for 100uA, as opposed to ~10000 in 
    LO gain)